5/36f = 5/16
6 8/21h = 13 1/3
5d = 3/10 This is algebra and im really confused.

Answer by rfer(16322)

A)
5/36f=5/16
f=(5/16)/(5/36)
f=5/16*36/5
f=180/80
f=2 1/4
B)
6 8/21h=13 1/3
h=(13 1/3)/(6 8/21)
h=40/3*21/134
h=840/402
h=2 6/67
C)
5d=3/10
d=(3/10)/(5/1)
d=3/10*1/5
d=3/50